What if the equations are 7r-6s=42 and 6r+7s=75 ? Would you multiply both equations by -6?I have done it like that I get an answer that I do not think is right.

Answer by checkley77(12569)

7r-6s=42 if you're going to use the elimination method then you need to multipy either the s term in the first equation by 7 & then s term in the second equation by 6 & then add them.
6r+7s=75
************************************
49r-42s=294
36r-42s=450
-----------------
85r=744
r=744/85 or 8.75294 ans.
7*8.75294-6s=42
61.2706-6s=42
-6x=42-61.2706
-6x=-19.27056
x=-19.27056/-6
x=3.2118 ans
Proof:
6*8.75294+7*3.2118=75
52.51764+22.4826=75
75~75